# On the Use of Hierarchical Vision Foundation Models for Low-Cost Human Mesh Recovery and Pose Estimation

### v1

@@ Abstract
In this work, we aim to develop simple and efficient models for human mesh recovery (HMR) and its predecessor task, human pose estimation (HPE). State-of-the-art HMR methods, such as HMR2.0 and its successors, rely on large, non-hierarchical vision transformers as encoders, which are inherited from the corresponding HPE models like ViTPose. To establish baselines across varying computational budgets, we first construct three lightweight HMR2.0 variants by adapting the corresponding ViTPose models. In addition, we propose leveraging the early stages of hierarchical vision foundation models (VFMs), including Swin Transformer, GroupMixFormer, and VMamba, as encoders. This design is motivated by the observation that intermediate stages of hierarchical VFMs produce feature maps with resolutions comparable to or higher than those of non-hierarchical counterparts. We conduct a comprehensive evaluation of 27 hierarchical-VFM-based HMR and HPE models, demonstrating that using only the first two or three stages achieves performance on par with full-stage models. Moreover, we show that the resulting truncated models exhibit better trade-offs between accuracy and computational efficiency compared to existing lightweight alternatives.
@@
Figure 1: (Left): 3D pose estimation results (PA-MPJPE; lower is better) of human mesh recovery (HMR) models on the Human3.6M dataset [29]. (Right): 2D pose estimation results (AP; higher is better) of human pose estimation (HPE) models on the COCO-val dataset [53]. Circle size reflects model size. Gray circles indicate existing methods, while colored circles represent our proposed models, which leverage the early stages of hierarchical vision foundation models (VFMs) as encoders. These results demonstrate that our models tend to offer a more favorable trade-off between performance and computational cost (i.e., GFLOPs) compared to existing approaches.
@@ 1 Introduction
Human mesh recovery (HMR) plays a central role in a wide range of applications, including animation, virtual try-on, sports analytics, and human-computer interaction [118, 85, 23, 54]. Over the past decade, this research field has witnessed remarkable progress, driven in part by vision foundation models (VFMs) [28, 88, 78, 15]. While early HMR approaches [32, 112, 39, 47, 37] primarily relied on convolutional neural networks (CNNs), recent state-of-the-art (SoTA) methods [11, 16, 51, 52, 1, 62, 9, 69, 50, 20] have increasingly adopted Transformer-based architectures [87]. Among them, HMR2.0 [26] has garnered significant attention for its simplicity and strong performance. HMR2.0 and its successors [63, 22, 70, 66, 75, 17] employ a large non-hierarchical vision transformer (i.e., ViT-H [15]) as their encoder, which is inherited from the corresponding human pose estimation (HPE) model (i.e., ViTPose-H [102]).
@@
In general, even for HMR and HPE, large VFMs demand substantial computational resources, which can hinder their deployment in real-time or resource-constrained settings such as mobile devices or edge computing environments. With HMR2.0, a straightforward approach to alleviate this issue is to use smaller ViT variants (e.g., ViT-L, ViT-B, ViT-S) as encoders. Since this direction has not been explored in the literature, in this work we instantiate smaller variants of HMR2.0 as baselines (see §3.2.1 for details). Beyond this, to better balance performance and efficiency while preserving the architectural simplicity of HMR2.0, we investigate the use of hierarchical VFMs [56, 24, 55] as encoders for HMR and its predecessor, HPE. The key insight motivating our approach lies in the resolution characteristics of intermediate representations in hierarchical VFMs. They typically follow a four-stage structure, where the spatial resolution of feature maps are higher or the same with the consistent resolution seen in non-hierarchical VFMs. Therefore, if the intermediate outputs of pretrained hierarchical VFMs retain sufficient semantic richness and spatial detail, the latter stages of the original backbone can be removed. This allows for reductions in model size and computational cost without compromising architectural simplicity.
@@
We conduct extensive experiments to validate this observation by instantiating HMR and HPE models within the HMR2.0 and ViTPose frameworks, using different stages of three hierarchical VFMs as encoders: Swin Transformer [56], GroupMixFormer [25], and VMamba [55]. In total, we instantiate 27 hierarchical-VFM-based HMR and HPE models. Our results consistently show that models using only the first two or three VFM stages as encoders achieve performance comparable to, and occasionally better than, their full four-stage counterparts. Moreover, these truncated models demonstrate a more favorable trade-off between accuracy and efficiency compared to existing lightweight approaches, including the small ViT-based variants of HMR2.0 and ViTPose. Our contributions are summarized as follows:
@@
  • We investigate the use of hierarchical VFMs [56, 24, 55] as encoders within the HMR2.0 and ViTPose frameworks. Experimental results show that models utilizing only the first few VFM stages as their encoders achieve performance comparable to, and sometimes better than, their full four-stage counterparts for both HMR and HPE.
@@ 2.1 Human Mesh Recovery (HMR)
HMR has been extensively studied under a variety of problem settings, including multi-person HMR [69, 48, 81, 82, 77, 5, 79, 13, 110, 111, 67], multi-view HMR [7, 114, 119, 46, 65, 109, 30, 100], video-based HMR [116, 12, 36, 94, 34, 74, 106], full-body HMR [64, 91, 72, 113, 42, 49, 8], privacy-preserving HMR [23, 97, 25], and prompt-based HMR [92, 50, 20]. In this work, we focus on the most fundamental setting, image-based HMR, where the objective is to predict SMPL [58] parameters from a single cropped image of a person.
@@
While early works employed CNNs such as ResNet [28], HRNet [78, 88], and EfficientNet [83] as backbone encoders [21, 40, 60, 32, 112, 39, 73, 71, 98, 47, 37, 43, 103], they have been superseded by Transformer-based methods in recent years. For instance, METRO [51, 11], Mesh Graphormer [52] and PointHMR [35] introduce hybrid CNN-Transformer architectures that leverage pretrained CNN features while capturing global dependencies via self-attention. On the other hand, several recent works [26, 49, 9, 8] utilize fully Transformer-based encoders. HMR2.0 [26], along with its successors [63, 22, 70, 66, 75, 17], employs ViT-H as the encoder, initialized through pretraining on HPE tasks. SMPLer-X [8], uses four ViTs to construct full-body HMR models with varying model sizes. More recently, prompt-based HMR approaches including ChatPose [20] and ChatHuman [50] leverage the ViT encoder from CLIP [68] to align visual and textual features.
@@
Previous works have primarily explored the use of non-hierarchical transformers, ViT or DeiT, for the HMR task. In contrast, this study investigates the application of hierarchical VFMs, encompassing not only transformer-based architectures but also recently proposed state space models (SSMs). A notable exception in the literature is [62], which integrates hierarchical transformers such as the Swin Transformer [56, 57] and Twins [14] into the HMR framework of [32]. DeFormer [107] also explores the use of hierarchical transformers, such as the Swin Transformer with a Feature Pyramid Network and the Mix Transformer [10, 99], as backbones. Nevertheless, as discussed in §1, our work goes beyond simply applying hierarchical VFMs to HMR: We further investigate the use of only the initial stages of these hierarchical models as encoders, aiming to develop more efficient HMR architectures.
@@
Several studies have aimed to develop efficient HMR models [117, 16, 1, 93]. For instance, CoarseMETRO [1] employs a coarse-to-fine strategy to reduce the computational burden of early transformer layers, while TORE [16] accelerates HMR by pruning background tokens. POTTER [117] integrates a high-resolution stream with a basic stream to recover more accurate human meshes while reducing memory usage and computational cost. Although effective, these methods often introduce additional architectural complexity. In contrast, our approach, leveraging the early stages of hierarchical VFMs, offers a simpler alternative that is potentially complementary to these techniques.
@@ 2.2 Human Pose Estimation (HPE)
Since the encoders in HMR methods are often inherited from those used in HPE models, another promising direction for developing efficient HMR is to adapt efficient HPE approaches. Several studies have explored this in the context of HPE. For example, MEMe [41], Lite-HRNet [108], and LitePose [91] focus on optimizing popular CNN backbones such as EfficientNet [83] and HRNet [78, 88] to improve the trade-off between accuracy and efficiency. DANet [59] proposes an improved multi-scale feature fusion strategy that eliminates the need for computationally expensive cascaded pyramid architectures. SimCC [45] reformulates HPE as two independent classification tasks for horizontal and vertical coordinates. Additionally, CNF [105] and ViPNAS [101] apply neural architecture search to automatically optimize network structures for improved efficiency.
@@ 2.3 Vision Foundation Model (VFM)
VFMs can be broadly categorized into non-hierarchical models [15, 86, 76, 6, 19, 84, 95, 104, 2, 105] and hierarchical models [56, 18, 89, 112, 44, 90, 80, 115, 24, 55]. Following the discussion above, this work focuses on recent hierarchical VFMs. Among them, we select Swin Transformer (Swin) [56], GroupMixFormer (GMF), and VMamba (VM) [57] based on their strong performance, widespread recognition, and the availability of public code with pretrained weights. Please refer to the original papers for detailed architectural descriptions. All three models consist of four VFM stages, with output feature map resolutions of $1/4\times 1/4$, $8\times 1/8, 1/1$$16\times 1/16$, and $1/32\times 1/32$ relative to the input image resolution at stages 1 through 4, respectively. In §4, we will explore to use their first few stages as HMR and HPE encoders.
HMR2.0 [26] HMR2.0-L HMR2.0-B HMR2.0-S
@@
Table 1: Building on the original HMR2.0 architecture proposed by [26], we introduce three scaled variants: HMR2.0-L, HMR2.0-B, and HMR2.0-S. NN denotes the number of Transformer layers, $h$ represents the number of attention heads in each cross-attention layer, $d_{\text{hid}}$ is the hidden dimension size, and $d_{\text{ff}}$ refers to the hidden dimension size of the feed-forward MLP block.
@@ 3 Baseline for HMR & HPE
To ensure better self-containment, we briefly review ViTPose [102] and HMR2.0 [26] in this section, as they serve as baselines for HPE and HMR, respectively. Both ViTPose and HMR2.0 employ a non-hierarchical VFM, ViT [15] as their encoder. While ViTPose explores four ViT variants of different sizes, HMR2.0 utilizes only the largest variant. To establish a more comprehensive baseline, we introduce smaller ViT-based variants of HMR2.0 in § 3.2.1, using encoders inherited from the corresponding ViTPose models.
3.1 ViTPose [102]
As illustrated in Figure 2 (a), ViTPose adopts a straightforward encoder-decoder architecture: a ViT serves as the encoder to generate a feature map, while the decoder comprises deconvolution layers followed by a prediction layer that outputs heatmaps corresponding to keypoints. Given an image of a person with height $H$ and width $W$ (typically, $H=256$ and $W=192$), the encoder produces features with a spatial resolution of $H/16\times W/16$. In this work, we utilize the official ViTPose repository [^1] https://github.com/ViTAE-Transformer/ViTPose, including pretrained weights, to deploy four model variants, i.e., ViTPose-H, ViTPose-L, ViTPose-B, and ViTPose-S, all of which are finetuned on the COCO dataset [53]. The results, including model size, computational complexity (GFLOPs), and frames per second (FPS), are summarized in Table 3.
3.2 HMR2.0 [26]
HMR2.0 also adopts an encoder-decoder architecture. As illustrated in Figure 2 (b), a ViT-based encoder produces a feature map, which is subsequently flattened and passed to a transformer-based decoder to predict the SMPL [58] parameters, i.e., the pose parameters $\alpha$, shape parameters $\beta$, and camera parameters $\theta$. During training, the encoder is initialized with pretrained weights from the ViTPose. Note that in [26], only ViTPose-H is used as the encoder. Its architectural specifications are provided in Table 1, and the corresponding 2D and 3D pose estimation results obtained using the official code [^2] https://github.com/shubham-goel/4D-Humans are reported in Table 4.

### v2

@@ Abstract
In this work, we aim to develop simple and efficient models for human mesh recovery (HMR) and its predecessor task, human pose estimation (HPE). State-of-the-art HMR methods, such as HMR2.0 and its successors, rely on large, non-hierarchical vision transformers as encoders, which are inherited from the corresponding HPE models like ViTPose. To establish baselines across varying computational budgets, we first construct three lightweight HMR2.0 variants by adapting the corresponding ViTPose models. In addition, we propose leveraging the early stages of hierarchical vision foundation models (VFMs), including Swin Transformer, GroupMixFormer, and VMamba, as encoders. This design is motivated by the observation that intermediate stages of hierarchical VFMs produce feature maps with resolutions comparable to or higher than those of non-hierarchical counterparts. We conduct a comprehensive evaluation of 27 hierarchical-VFM-based HMR and HPE models, demonstrating that using only the first two or three stages achieves performance on par with full-stage models. Moreover, we show that the resulting truncated models exhibit better trade-offs between accuracy and computational efficiency compared to existing lightweight alternatives. The source code is available at https://github.com/nttcom/TruncHierVFM.
@@
Figure 1: (Left): 3D pose estimation results (PA-MPJPE; lower is better) of human mesh recovery (HMR) models on the Human3.6M dataset [ionescu+2014tpami]. (Right): 2D pose estimation results (AP; higher is better) of human pose estimation (HPE) models on the COCO-val dataset [lin+2014eccv]. Circle size reflects model size. Gray circles indicate existing methods, while colored circles represent our proposed models, which leverage the early stages of hierarchical vision foundation models (VFMs) as encoders. These results demonstrate that our models tend to offer a more favorable trade-off between performance and computational cost (i.e., GFLOPs) compared to existing approaches.
@@ 1 Introduction
Human mesh recovery (HMR) plays a central role in a wide range of applications, including animation, virtual try-on, sports analytics, and human-computer interaction [zheng+2023acm, tian+2023tpami, gao+2025air, liu+2024neuro]. Over the past decade, this research field has witnessed remarkable progress, driven in part by vision foundation models (VFMs) [he+2016cvpr, wang+2021tpami, sun+2019cvpr, dosovitskiy+2020iclr]. While early HMR approaches [kanazawa+2018cvpr, zhang+2021iccv, kolotouros+2021iccv, li+2022eccv2, kocabas+2021iccv] primarily relied on convolutional neural networks (CNNs), recent state-of-the-art (SoTA) methods [cho+2022eccv, dou+2023iccv, lin+2021cvpr, lin+2021iccv, agarwal+2024eccvw, pang+2022neurips, cai+2024tpami, romain+2025cvprw, lin+2025cvpr, feng+2024cvpr] have increasingly adopted Transformer-based architectures [vaswani+2017nips]. Among them, HMR2.0 [goel+2023iccv] has garnered significant attention for its simplicity and strong performance. HMR2.0 and its successors [patel+20243dv, fiche+2025cvpr, saleem+2025aaai, prospero+2025cvprw, stathopoulos+2024cvpr, dwivedi+2024cvpr] employ a large non-hierarchical vision transformer (i.e., ViT-H [dosovitskiy+2020iclr]) as their encoder, which is inherited from the corresponding human pose estimation (HPE) model (i.e., ViTPose-H [xu+2022neurips]).
@@
In general, even for HMR and HPE, large VFMs demand substantial computational resources, which can hinder their deployment in real-time or resource-constrained settings such as mobile devices or edge computing environments. With HMR2.0, a straightforward approach to alleviate this issue is to use smaller ViT variants (e.g., ViT-L, ViT-B, ViT-S) as encoders. Since this direction has not been explored in the literature, in this work we instantiate smaller variants of HMR2.0 as baselines (see §3.2.1 for details). Beyond this, to better balance performance and efficiency while preserving the architectural simplicity of HMR2.0, we investigate the use of hierarchical VFMs [liu+2021iccv, ge+2023arxiv, liu+2024neurips] as encoders for HMR and its predecessor, HPE. The key insight motivating our approach lies in the resolution characteristics of intermediate representations in hierarchical VFMs. They typically follow a four-stage structure, where the spatial resolution of feature maps are higher or the same with the consistent resolution seen in non-hierarchical VFMs. Therefore, if the intermediate outputs of pretrained hierarchical VFMs retain sufficient semantic richness and spatial detail, the latter stages of the original backbone can be removed. This allows for reductions in model size and computational cost without compromising architectural simplicity.
@@
We conduct extensive experiments to validate this observation by instantiating HMR and HPE models within the HMR2.0 and ViTPose frameworks, using different stages of three hierarchical VFMs as encoders: Swin Transformer [liu+2021iccv], GroupMixFormer [ge+2024cvpr], and VMamba [liu+2024neurips]. In total, we instantiate 27 hierarchical-VFM-based HMR and HPE models. Our results consistently show that models using only the first two or three VFM stages as encoders achieve performance comparable to, and occasionally better than, their full four-stage counterparts. Moreover, these truncated models demonstrate a more favorable trade-off between accuracy and efficiency compared to existing lightweight approaches, including the small ViT-based variants of HMR2.0 and ViTPose. Our contributions are summarized as follows:
@@
  • We investigate the use of hierarchical VFMs [liu+2021iccv, ge+2023arxiv, liu+2024neurips] as encoders within the HMR2.0 and ViTPose frameworks. Experimental results show that models utilizing only the first few VFM stages as their encoders achieve performance comparable to, and sometimes better than, their full four-stage counterparts for both HMR and HPE.
@@ 2.1 Human Mesh Recovery (HMR)
HMR has been extensively studied under a variety of problem settings, including multi-person HMR [romain+2025cvprw, liao+2024tvcg, sun+2021iccv, sun+2022cvpr, su+2025cvpr, fabien+2024eccv, sun+2024cvpr, choi+2022cvpr, zanfir+2018cvpr, zanfir+2018neurips, qiu+2023cvpr], multi-view HMR [bogo+eccv2016, zhang+2023iccv, zhu+2025aaai, li+2021wacv, pavlakos+2022cvpr, yu+2022neurips, jia+2023aaai, xie+2024icmr], video-based HMR [zhang+2025cvpr, choi+2021cvpr, kocabas+2020cvpr, wei+2022cvpr, kang+2025cvprw, shen+2024sa, yang+2023iclr], full-body HMR [pavlakos+2019cvpr, wang+2022cvpr, shen+2024acmmm, zhang+2023tpami, li+2025tpami, lin+2023cvpr, cai+2023neurips], privacy-preserving HMR [gao+2025air, wu+2025cvpr, ge+2024cvpr], and prompt-based HMR [wang+2025cvpr, lin+2025cvpr, feng+2024cvpr]. In this work, we focus on the most fundamental setting, image-based HMR, where the objective is to predict SMPL [loper+2015tog] parameters from a single cropped image of a person.
@@
While early works employed CNNs such as ResNet [he+2016cvpr], HRNet [sun+2019cvpr, wang+2021tpami], and EfficientNet [tan+2019icml] as backbone encoders [fiche+2024eccv, le+2024cvpr, ma+2023cvpr, kanazawa+2018cvpr, zhang+2021iccv, kolotouros+2021iccv, shen+2025icml, sarandi+2024neurips, xiao+2024eccv, li+2022eccv2, kocabas+2021iccv, li+2024tip, xu+2024cvpr], they have been superseded by Transformer-based methods in recent years. For instance, METRO [lin+2021cvpr, cho+2022eccv], Mesh Graphormer [lin+2021iccv] and PointHMR [kim+2023cvpr] introduce hybrid CNN-Transformer architectures that leverage pretrained CNN features while capturing global dependencies via self-attention. On the other hand, several recent works [goel+2023iccv, lin+2023cvpr, cai+2024tpami, cai+2023neurips] utilize fully Transformer-based encoders. HMR2.0 [goel+2023iccv], along with its successors [patel+20243dv, fiche+2025cvpr, saleem+2025aaai, prospero+2025cvprw, stathopoulos+2024cvpr, dwivedi+2024cvpr], employs ViT-H as the encoder, initialized through pretraining on HPE tasks. SMPLer-X [cai+2023neurips], uses four ViTs to construct full-body HMR models with varying model sizes. More recently, prompt-based HMR approaches including ChatPose [feng+2024cvpr] and ChatHuman [lin+2025cvpr] leverage the ViT encoder from CLIP [radford+2021icml] to align visual and textual features.
@@
Previous works have primarily explored the use of non-hierarchical transformers, ViT or DeiT, for the HMR task. In contrast, this study investigates the application of hierarchical VFMs, encompassing not only transformer-based architectures but also recently proposed state space models (SSMs). A notable exception in the literature is [pang+2022neurips], which integrates hierarchical transformers such as the Swin Transformer [liu+2021iccv, liu+2022cvpr] and Twins [chu+2021neurips] into the HMR framework of [kanazawa+2018cvpr]. DeFormer [yoshiyasu+2023cvpr] also explores the use of hierarchical transformers, such as the Swin Transformer with a Feature Pyramid Network and the Mix Transformer [cao+2022ijcai, xie+2021neurips], as backbones. Nevertheless, as discussed in §1, our work goes beyond simply applying hierarchical VFMs to HMR: We further investigate the use of only the initial stages of these hierarchical models as encoders, aiming to develop more efficient HMR architectures.
@@
Several studies have aimed to develop efficient HMR models [zheng+2023cvpr, dou+2023iccv, agarwal+2024eccvw, wang+2025visapp]. For instance, CoarseMETRO [agarwal+2024eccvw] employs a coarse-to-fine strategy to reduce the computational burden of early transformer layers, while TORE [dou+2023iccv] accelerates HMR by pruning background tokens. POTTER [zheng+2023cvpr] integrates a high-resolution stream with a basic stream to recover more accurate human meshes while reducing memory usage and computational cost. Although effective, these methods often introduce additional architectural complexity. In contrast, our approach, leveraging the early stages of hierarchical VFMs, offers a simpler alternative that is potentially complementary to these techniques.
@@ 2.2 Human Pose Estimation (HPE)
Since the encoders in HMR methods are often inherited from those used in HPE models, another promising direction for developing efficient HMR is to adapt efficient HPE approaches. Several studies have explored this in the context of HPE. For example, MEMe [li+2022sensors], Lite-HRNet [yu+2021cvpr], and LitePose [wang+2022cvpr] focus on optimizing popular CNN backbones such as EfficientNet [tan+2019icml] and HRNet [sun+2019cvpr, wang+2021tpami] to improve the trade-off between accuracy and efficiency. DANet [luo+2021icme] proposes an improved multi-scale feature fusion strategy that eliminates the need for computationally expensive cascaded pyramid architectures. SimCC [li+2022eccv] reformulates HPE as two independent classification tasks for horizontal and vertical coordinates. Additionally, CNF [yang+2022pr] and ViPNAS [xu+2021cvpr] apply neural architecture search to automatically optimize network structures for improved efficiency.
@@ 2.3 Vision Foundation Model (VFM)
VFMs can be broadly categorized into non-hierarchical models [dosovitskiy+2020iclr, touvron+2021icml, steiner+2022tmlr, beyer+2023cvpr, fang+2024ivc, tian+2023cvpr, wei+2023iccv, yang+2023iclr2, alkin+2025iclr, yang+2022pr] and hierarchical models [liu+2021iccv, fan+2021iccv, wang+2021iccv, zhang+2021iccv, li+2022cvpr, wang+2022cvmj, sun+2023tpami, zhang+2023iclr, ge+2023arxiv, liu+2024neurips]. Following the discussion above, this work focuses on recent hierarchical VFMs. Among them, we select Swin Transformer (Swin) [liu+2021iccv], GroupMixFormer (GMF), and VMamba (VM) [liu+2022cvpr] based on their strong performance, widespread recognition, and the availability of public code with pretrained weights. Please refer to the original papers for detailed architectural descriptions. All three models consist of four VFM stages, with output feature map resolutions of $1/4\times 1/4$, $8\times 1/8, 1/1$$16\times 1/16$, and $1/32\times 1/32$ relative to the input image resolution at stages 1 through 4, respectively. In §4, we will explore to use their first few stages as HMR and HPE encoders.
HMR2.0 [goel+2023iccv] HMR2.0-L HMR2.0-B HMR2.0-S
@@
Table 1: Building on the original HMR2.0 architecture proposed by [goel+2023iccv], we introduce three scaled variants: HMR2.0-L, HMR2.0-B, and HMR2.0-S. NN denotes the number of Transformer layers, $h$ represents the number of attention heads in each cross-attention layer, $d_{\text{hid}}$ is the hidden dimension size, and $d_{\text{ff}}$ refers to the hidden dimension size of the feed-forward MLP block.
@@ 3 Baseline for HMR & HPE
To ensure better self-containment, we briefly review ViTPose [xu+2022neurips] and HMR2.0 [goel+2023iccv] in this section, as they serve as baselines for HPE and HMR, respectively. Both ViTPose and HMR2.0 employ a non-hierarchical VFM, ViT [dosovitskiy+2020iclr] as their encoder. While ViTPose explores four ViT variants of different sizes, HMR2.0 utilizes only the largest variant. To establish a more comprehensive baseline, we introduce smaller ViT-based variants of HMR2.0 in § 3.2.1, using encoders inherited from the corresponding ViTPose models.
3.1 ViTPose [xu+2022neurips]
As illustrated in Figure 2 (a), ViTPose adopts a straightforward encoder-decoder architecture: a ViT serves as the encoder to generate a feature map, while the decoder comprises deconvolution layers followed by a prediction layer that outputs heatmaps corresponding to keypoints. Given an image of a person with height $H$ and width $W$ (typically, $H=256$ and $W=192$), the encoder produces features with a spatial resolution of $H/16\times W/16$. In this work, we utilize the official ViTPose repository [^1] https://github.com/ViTAE-Transformer/ViTPose, including pretrained weights, to deploy four model variants, i.e., ViTPose-H, ViTPose-L, ViTPose-B, and ViTPose-S, all of which are finetuned on the COCO dataset [lin+2014eccv]. The results, including model size, computational complexity (GFLOPs), and frames per second (FPS), are summarized in Table 3.
3.2 HMR2.0 [goel+2023iccv]
HMR2.0 also adopts an encoder-decoder architecture. As illustrated in Figure 2 (b), a ViT-based encoder produces a feature map, which is subsequently flattened and passed to a transformer-based decoder to predict the SMPL [loper+2015tog] parameters, i.e., the pose parameters $\alpha$, shape parameters $\beta$, and camera parameters $\theta$. During training, the encoder is initialized with pretrained weights from the ViTPose. Note that in [goel+2023iccv], only ViTPose-H is used as the encoder. Its architectural specifications are provided in Table 1, and the corresponding 2D and 3D pose estimation results obtained using the official code [^2] https://github.com/shubham-goel/4D-Humans are reported in Table 4.